# Tao Yu,  Zongdian Li,  Omar Hashash,  Kei Sakaguchi,  Walid Saad,  and Mérouane Debbah

### v1

@@ III-B1 Real-time computing for data collection and actuation
Real-time computing in the IoFDT is essential for effectively coordinating multiple DTs with one another and with their dynamic physical systems. This requires not only efficient architectural design, but agile resource allocation strategies for rapid processing of complex datasets on distributed edge servers, which often have limitations in computing power. The challenges in real-time computing are multifaceted. Latency issues arise from data transmission and processing across distributed computational nodes, and any delay can degrade the decision-making. Edge servers must process large-scale data inflows, potentially causing bottlenecks. The computational infrastructure and resources require high flexibility, scaling to meet DTs’ real-time needs. Consider smart mobility services in which autonomous vehicles’ safety and efficiency hinge on immediate sensor data processing and communication with surrounding infrastructures. Any lag in these processes could result in outdated decisions, risking safety and operational inefficiencies. Moreover, when multiple DTs interact, the challenge escalates as each DT requires data from others in real-time to make informed decisions. This interdependence necessitates not only rapid data sharing but synchronized data processing and analysis across DTs, which demands an advanced level of computational orchestration for system-wide coherence and efficiency. To overcome these challenges, a potential solution is implementing edge computing architectures with specialized accelerators for data-intensive tasks, such as FPGAs and GPUs, which can significantly enhance real-time analytics processing at edges. Additionally, employing predictive resource allocation strategies utilizing machine learning models to forecast demand and preemptively adjust resource distribution can improve DTs’ responsiveness. This predictive approach, combined with flexible edge computing resources, would minimize latency and enable more effective real-time computation and actuation for tasks that are critical to operation and synchronization of DTs, such as immediate traffic re-routing in connected vehicle networks.
@@ III-B2 Scalability
The scalability of the IoFDT infrastructure is a key challenge, because the IoFDT encompasses diverse, geographically dispersed DTs representing increasingly complex and dynamic physical systems. As the number and complexity of DTs grow, computational demands and data volumes in IoFDT also grow, necessitating a distributed computing network capable of dynamic adjustment and scaling. Key scalability challenges include developing a computing architecture ensuring high-throughput data processing and rapid elasticity for growing and fluctuating workloads, while also being resilient to node failures to maintain operational integrity and system performance. A potential solution to these challenges lies in advanced distributed computing networks with dynamic adaptive load balancing, auto-scaling, and resource provisioning responding to DTs’ increasing computational demands. Furthermore, spectral graph theory offers a powerful tool for modeling the complex network of DT interconnections. By treating the network as a graph, spectral analysis enables us to identify the interactivity degrees of DTs and computational nodes in IoFDT. Such insight allows for optimized resource allocation, and thereby can optimize network performance and reduce computational overhead. Essentially, spectral graph theory also aids streamlining communication routes and prioritizing resource distribution to enhance the overall efficiency and reduce latency in IoFDT.
@@ III-C1 AI Models for IoFDT Design
The IoFDT requires new AI frameworks for interconnected and dynamic DTs evolving with data and real-time actuation. At the core, each DT is an AI model that continuously updates its physical space twin, and this is where AI is leveraged to create DTs and the entire IoFDT system. Learning within IoFDT must consider networking constraints while accurately twinning distributed physical processes across various dimensions. To fulfill network-aware AI twinning, AI must be adept at creating context-aware twin models that adapt with data inflows, and extract complex interrelations across diverse, heterogeneous systems. These frameworks must encapsulate the capability for multidimensional representations that accurately reflect physical processes and inform network design within system-wide communication and computing limitations. Moreover, compared to standalone DTs, generating and leveraging synthetic data for IoFDT to simulate hypothetical yet plausible scenarios pose greater challenges, as clusters of federated DTs involve increased complexity of dynamic, diverse, interconnected systems. The adoption of continual graph neural networks (CGNNs)[10], combining benefits of continual learning (CL) with graph neural networks (GNNs)[11], promises to address these demands. On the one hand, CL provides an agile incremental learning method with swift model updates [12]. On the other hand, GNNs present an effective method to model the interrelations between the multi-level DT elements. Hence, CGNNs can adapt to temporal dynamics and complex interactions inherent in IoFDT, enhancing the resilience and adaptability of the network.
@@ IV Practical Implementation and Realization of IoFDT: Platform and Experimental Field
In addition to theoretical challenges outlined in Section III, a unified, experimental IoFDT platform is needed to blend all technical elements and orchestrates their interactions. Indeed, experiments with practical IoFDT services on a fully functional experimental platform are indispensable to validate fundamental research and generate testing data for system design refinement.
@@ IV-A Unified IoFDT Platform to Integrate All Elements
We propose an IoFDT platform, as a universal integrator and orchestrator, to unify all elements in IoFDT for seamless interactions. To realize such a platform, three functional modules become essential: the DT orchestrator, the network orchestrator, and the computing orchestrator.
@@
DT Orchestrator: This orchestrator manages DTs throughout life-cycle, including storage, creation, distribution, operation, and interaction. Central to this module are repositories of reusable DT functions and knowledge bases (AI models). For new applications and associated DTs, instead of starting from the ground up, they leverage pre-existing function modules and knowledge, accelerating DT development and implementation. The DT orchestrator distributes and deploys new DTs to optimized edge or cloud servers based on application scenarios and QoS requirements. It orchestrates DT interactions in the IoFDT, including data sharing and learning through DT federalization, data format/protocols coordination, data desensitization, and knowledge base management. In this containerization approach, the orchestrator assembles components to create DTs. These components, along with function modules, dependencies, configurations, AI models, and desensitized data, are then packaged into containers using tools such as Docker and containerd. They are then deployed and run by tools such as Kubernetes and Docker Swarm. The DT orchestrator, with a global system view, also decides the federation participation of new DTs, considering function, service, and resources. A continuous integration/continuous deployment (CI/CD) pipeline using automation tools like Jenkins ensures this process for each DT creation or update. Middle-ware is also necessary for data interoperability between networks, DTs, and devices using different data formats/protocols.
@@
Network Orchestrator: This orchestrator manages the data flow and resource in IoFDT by allocating network resources, managing diverse communication protocols, and maintaining network QoS for each DT to collect data, process it, and send the input back to the physical space in real-time. Software-defined networking (SDN) and network function virtualization (NFV) can be employed for dynamic and programmable networks through an IoFDT network control plane. To orchestrate IoFDT networks and enable access to customized virtual network functions (VNFs) for each DT, SDN controllers provide a global, real-time, and unified view of resources and programmable interfaces to manipulate data flows. In practical implementation, SDN frameworks, such as OpenDaylight, Open Network Operating System, and Ryu, can be leveraged to abstract resources in networks, supporting the dynamic slicing of networks for diverse DTs. RAN and core network functions are transformed into VNFs and managed by open platforms, such as Open RAN (O-RAN). An interface between the chosen SDN controller and the VNF management platform is necessary to monitor the status of VNFs. By implementing this composition, the IoFDT network orchestrator maps the resources required by DT to the specific physical infrastructures, controlling data flows to ensure seamless connectivity.
@@
Computing Orchestrator: The computing orchestrator manages computing resources, both cloud and edge, to ensure that each DT can process the data and learn the physical space in real-time. It involves such as resource monitoring, workload balancing, dynamic resource allocation, and edge-cloud coordination. Such orchestration must be in real-time while constantly adapting to the dynamic needs of the DTs and the fluctuating availability of resources. Practically, each DT application can be packaged into a unit, such as a Pod in Kubernetes comprising DT and knowledge base containers. It can be deployed in a cloud-edge server computing cluster. The orchestrator, typically at the cloud server, monitors real-time resource usage in the cluster and DT demands, including memory, CPU, GPU, and storage, as assessed by tools such as Kubelet. With the global insights and perspective, the orchestrator optimally allocates computing resources, ensuring real-time processing capabilities for DTs.
@@ IV-B Experimental Platform for IoFDT
Fig. 4 shows a proposed architecture of an IoFDT PoC implementation with physical systems, network infrastructure, and DTs. It is a open, programmable, and universal platform, constructed with state-of-the-art, programmable sensing, communication, and computing hardware (e.g., LiDARs, 5G NR networks/ITS networks, MEC/cloud servers), and developer-friendly open-source software (e.g., Open Daylight, Robot Operating System, and TensorFlow/PyTorch). The experimental integrated IoFDT platform unifies cyber and physical spaces and satisfy the practical requirements of diverse applications. Given the platform’s emphasis on mobility as an example, we developed a Smart Mobility Research & Education Field in Tokyo Tech, as shown in Fig. 4, with state-of-the-art infrastructures, e.g., 5G NR network, millimeter (mmWave) vehicular networks (backhaul/access)[14], roadside units (RSUs) and connected automated driving vehicles (CAVs) as a cornerstone of PoC implementation for IoFDT.
@@
PoC trials demonstrated the system functions and enhanced vehicle perception, assisting safer automated driving. For example, as shown in 5, we developed a smart mobility DT fusing detected traffic information from RSUs and CAVs [15]. Based on real-time global traffic information in DT, optimized paths are calculated for CAV at intersections in the cloud server and sent to CAV via vehicular network. CAV will adjust its path accordingly. A demo video is available [^1] https://youtu.be/Y3XXlNIRmXI. The IoFDT PoC platform will be built on these preliminary experimental works as well as extensive works on AI, communication networks, computing networks, and platforms discussed in previous sections. We are also developing and implementing diverse smart mobility prototype applications on the IoFDT PoC system, aiming to demonstrate IoFDT’s feasibility and merits. Examples include DT-based collision prediction and avoidance system, and a DT-based car-sharing system, as typical and key Society 5.0 use-cases. Additionally, this platform can be utilized to develop a wide range of DT-based applications beyond smart mobility.
@@ V Conclusion and Recommendations
  • Synchronization and Interaction: While individual DTs must be synchronized with physical system, in an IoFDT, it is necessary to also look at the cross-DT interactions and coordination.
  • Adaptive and Continuous Learning Frameworks: AI-native networks and learning frameworks that are capable of changing and adapting over time are needed to create to support the IoFDT. These frameworks must be responsive to new data and changes in physical systems to continuously enhance DTs’ performance and reliability.
  • IoFDT-centric Resource Management: Due to the specific requirements of each DT in the IoFDT ecosystem, efficient allocation of network and computing resources, across the entire federation of DTs, is crucial for the IoFDT. This includes the implementation of learning-based dynamic network slicing and context-aware resource optimization.
  • Application and Service: The development of novel applications and services enabled by IoFDT is necessary, and Society 5.0 offers a fertile ground for this.

### v2

@@ III-B1 Real-time computing for data collection and actuation
Real-time computing in the IoFDT is essential for effectively coordinating multiple DTs with one another and with their dynamic physical systems. This requires not only efficient architectural design, but agile resource allocation strategies for rapid processing of complex datasets on distributed edge servers, which often have limitations in computing power. The challenges in real-time computing are multifaceted. Latency issues arise from data transmission and processing across distributed computational nodes, and any delay can degrade decision-making. Edge servers must process large-scale data inflows, potentially causing bottlenecks. The computational infrastructure and resources require high flexibility, scaling to meet DTs’ real-time needs. Consider smart mobility services in which autonomous vehicles’ safety and efficiency hinge on immediate sensor data processing and communication with surrounding infrastructures. Any lag in these processes could result in outdated decisions, risking safety and operational inefficiencies. Moreover, when multiple DTs interact, the challenge escalates as each DT requires data from others in real-time to make informed decisions. This interdependence necessitates not only rapid data sharing but synchronized data processing and analysis across DTs, which demands an advanced level of computational orchestration for system-wide coherence and efficiency. To overcome these challenges, a potential solution is implementing edge computing architectures with specialized accelerators for data-intensive tasks, such as FPGAs and GPUs, which can significantly enhance real-time analytics processing at edges. Employing predictive resource allocation strategies utilizing machine learning models to forecast demand and preemptively adjust resource distribution can improve DTs’ responsiveness. This predictive approach, combined with flexible edge computing resources, would minimize latency and enable more effective real-time computation and actuation for tasks that are critical to DTs’ operation and synchronization, such as immediate traffic re-routing in connected vehicle networks.
@@ III-B2 Scalability
The scalability of the IoFDT infrastructure is a key challenge, because the IoFDT encompasses diverse, geographically dispersed DTs representing increasingly complex and dynamic physical systems. As DTs’ number and complexity grow, computational demands and data volumes in IoFDT also grow, necessitating a distributed computing network capable of dynamic adjustment and scaling. Key scalability challenges include developing a computing architecture ensuring high-throughput data processing and rapid elasticity for growing and fluctuating workloads, while also being resilient to node failures to maintain operational integrity and system performance. A potential solution to these challenges lies in advanced distributed computing networks with dynamic adaptive load balancing, auto-scaling, and resource provisioning responding to DTs’ increasing computational demands. Furthermore, spectral graph theory offers a powerful tool for modeling the complex network of DT interconnections. By treating the network as a graph, spectral analysis enables us to identify the interactivity degrees of DTs and computational nodes in IoFDT. Such insight allows for optimized resource allocation, and thereby can optimize network performance and reduce computational overhead. Essentially, spectral graph theory also aids in streamlining communication routes and prioritizing resource distribution to enhance overall efficiency and reduce latency in IoFDT.
@@ III-C1 AI Models for IoFDT Design
The IoFDT requires new AI frameworks for interconnected and dynamic DTs evolving with data and real-time actuation. At its core, each DT is an AI model that continuously updates its physical space twin, and this is where AI is leveraged to create DTs and the entire IoFDT system. Learning within IoFDT must consider networking constraints while accurately twinning distributed physical processes across various dimensions. To fulfill network-aware AI twinning, AI must be adept at creating context-aware twin models that adapt at data inflows, and extract complex interrelations across diverse, heterogeneous systems. These frameworks must encapsulate the capability for multidimensional representations that accurately reflect physical processes and inform network design within system-wide communication and computing limitations. Moreover, compared to standalone DTs, generating and leveraging synthetic data for IoFDT to simulate hypothetical yet plausible scenarios pose greater challenges, as clusters of federated DTs involve increased complexity of dynamic, diverse, interconnected systems. Continual graph neural network (CGNN)[11], combining benefits of continual learning (CL) with graph neural network (GNN), is a candidate promising to address these demands by adapting to temporal dynamics and complex interactions inherent in IoFDT. On the one hand, CL provides an agile incremental learning method with swift model updates [12]. On the other hand, GNNs present an effective method to model the interrelations between the multi-level DT elements. For instance, in smart manufacturing, IoFDT can integrate various DTs such as machinery, warehouse, and logistics. CGNNs are utilized to continuously learn and update relationships based on their complex data interactions and dependencies, thereby optimizing entire production process. When GNNs are not directly applicable, knowledge graphs can be employed to articulate and structure the non-graph-structured relationships.
@@ IV Practical Implementation and Realization of IoFDT: Platform and Experimental Field
In addition to theoretical challenges outlined in Section III, a unified, experimental IoFDT platform is needed to blend all technical elements and orchestrate their interactions. Indeed, experiments with practical Society 5.0 services in IoFDT on a fully functional experimental platform are indispensable to validate fundamental research and generate testing data for system design refinement.
@@ IV-A Unified IoFDT Platform to Integrate All Elements
We propose an IoFDT platform, as a universal integrator and orchestrator, to unify all elements in IoFDT and orchestrate seamless interactions among DTs. To realize such a platform, three functional modules become essential: the DT orchestrator, network orchestrator, and computing orchestrator.
@@
DT Orchestrator: This orchestrator manages DTs throughout life-cycle, including storage, creation, distribution, operation, and interaction. Central to this module are repositories of reusable DT functions and knowledge bases (AI models). For new applications and associated DTs, instead of starting from the ground up, they leverage pre-existing function modules and knowledge, accelerating DT development and implementation. The DT orchestrator distributes and deploys new DTs to optimized edge or cloud servers based on application scenarios and QoS requirements. It orchestrates DT interactions in the IoFDT, including data sharing and learning through DT federalization, data format/protocols coordination, data desensitization, and knowledge base management. In this containerization approach, the orchestrator assembles components to create DTs. These components, along with function modules, dependencies, configurations, AI models, and desensitized data, are then packaged into containers using tools such as Docker and containerd. They are then deployed and run by tools such as Kubernetes and Docker Swarm. The DT orchestrator, with a global system view, also decides the federation participation of new DTs, considering function, service, and resources. A continuous integration/continuous deployment (CI/CD) pipeline using automation tools like Jenkins ensures this process for each DT creation or update. Middleware is also necessary for data interoperability between networks, DTs, and devices using different data formats/protocols.
@@
Network Orchestrator: This orchestrator manages the data flow and resources in IoFDT by allocating network resources, managing diverse communication protocols, and maintaining network QoS for each DT to collect data, process it, and send the input back to the physical space in real-time. Software-defined networking (SDN) and network function virtualization (NFV) can be employed for dynamic and programmable networks through an IoFDT network control plane. To orchestrate IoFDT networks and enable access to customized virtual network functions (VNFs) for each DT, SDN controllers provide a global, real-time, and unified view of resources and programmable interfaces to manipulate data flows. In practical implementation, SDN frameworks, such as OpenDaylight, Open Network Operating System, and Ryu, can be leveraged to abstract resources in networks, supporting the dynamic slicing for diverse DTs. RAN and core network functions are transformed into VNFs and managed by open platforms, such as Open RAN (O-RAN). An interface between the chosen SDN controller and the VNF management platform is necessary to monitor the status of VNFs. By implementing this composition, the IoFDT network orchestrator maps the resources required by DT to the specific physical infrastructures, controlling data flows to ensure seamless connectivity.
@@
Computing Orchestrator: The computing orchestrator manages computing resources, both cloud and edge, to ensure that each DT can process the data and learn the physical space in real-time. It involves such as resource monitoring, workload balancing, dynamic resource allocation, and edge-cloud coordination. Such orchestration must be in real-time while constantly adapting to DTs’ dynamic needs and the fluctuating availability of resources. Practically, each DT application can be packaged into a unit, such as a Pod in Kubernetes comprising DT and knowledge base containers. It can be deployed in a cloud-edge server computing cluster. The orchestrator, typically at the cloud server, monitors real-time resource usage in the cluster and DT demands, including memory, CPU, GPU, and storage, as assessed by tools such as Kubelet. With global insights and perspective, the orchestrator optimally allocates computing resources, ensuring real-time processing capabilities for DTs.
@@ IV-B Experimental Platform for IoFDT
Fig. 4 shows a proposed architecture of an IoFDT PoC implementation with physical systems, network infrastructure, and DTs. It is an open, programmable, and universal platform, constructed with state-of-the-art, programmable sensing, communication, and computing hardware (e.g., LiDARs, 5G NR networks/ITS networks, MEC/cloud servers), and developer-friendly open-source software (e.g., Open Daylight, Robot Operating System, and TensorFlow/PyTorch). The experimental integrated IoFDT platform unifies cyber and physical spaces and satisfy the practical requirements of diverse applications. Given the platform’s emphasis on mobility as an example, we developed a Smart Mobility Research & Education Field in Tokyo Tech, as shown in Fig. 4, with state-of-the-art infrastructures, e.g., 5G NR network, millimeter (mmWave) vehicular networks (backhaul/access)[14], roadside units (RSUs) and connected automated driving vehicles (CAVs) as a cornerstone of PoC implementation for IoFDT.
@@
PoC trials demonstrated the system functions and enhanced vehicle perception, assisting safer automated driving. For example, as shown in Fig. 5, we have developed a smart mobility DT that fuses traffic information detected, processed, and then mapped into cyberspace at RSUs and CAVs[15]. This serves as an initial experimental implementation of integrating multiple local DTs to provide insights on the global traffic situation. A PoC application of path planning for CAVs was also conducted, in which optimized paths are calculated for CAVs based on real-time global traffic information in DT, and sent to CAV via vehicular network. The CAV then adjusts its path accordingly. A demo video of this PoC is available [^1] https://youtu.be/Y3XXlNIRmXI. The IoFDT PoC platform will be built on these preliminary experimental works as well as extensive works on AI, communication networks, computing networks, and platforms discussed in previous sections. We are also developing and implementing diverse smart mobility prototype applications on the IoFDT PoC system, aiming to demonstrate IoFDT’s feasibility and merits. Examples include DT-based collision prediction and avoidance system, and a DT-based car-sharing system, as typical and key Society 5.0 use-cases. Additionally, this platform can be utilized to develop a wide range of DT-based applications beyond smart mobility.
@@ V Conclusion and Recommendations
  • Synchronization and Interaction: While individual DTs must be synchronized with physical system, in an IoFDT, it is also necessary to look at cross-DT interactions and coordination.
  • Adaptive and Continuous Learning Frameworks: AI-native networks and learning frameworks that are capable of changing and adapting over time are needed to support the IoFDT. These frameworks must be responsive to new data and changes in physical systems to continuously enhance DTs’ performance and reliability.
  • IoFDT-centric Resource Management: Due to the specific requirements of each DT in the IoFDT ecosystem, efficient allocation of network and computing resources, across the entire DT federation, is crucial for the IoFDT. This includes the implementation of learning-based dynamic network slicing and context-aware resource optimization.
  • Privacy and Security: Given the interconnected nature of IoFDT, it is crucial to ensure privacy-preserving insights across DTs and maintain system integrity and security by, e.g., leveraging federated intelligence, robust encryption protocols, and access control mechanisms in the IoFDT framework.
  • Application and Service: The development of novel applications and services enabled by IoFDT is necessary, and Society 5.0 offers fertile ground for this.